# Solving the Poisson equation using coupled Markov chains

### v3

@@ B.6 Unbiased Poisson asymptotic variance estimator
Definition 51.
Let $H_{1}$ and $H_{2}$ be two independent unbiased estimators of $\pi(h)$ and $S$ be an unbiased estimator of $\pi(h^{2})$, all of the type described in Proposition 43. Let $R\in\{0,1,\ldots\}$ and $\Phi_{R}$ be as in Definition 49 and define
@@
Lemma 52.
Under Assumption 2, let $h\in L^{m}(\pi)$ for some $m>2\kappa/(\kappa-1)$. Let $H_{1}$, $H_{2}$ and $S$ be as in Definition 51. Then
@@
We note that if $h\in L^{m}(\pi)$ then $h^{2}\in L^{m/2}(\pi)$. Hence, we deduce by Proposition 43 that if $m/2>\kappa/(\kappa-1)$, i.e. $m>2\kappa/(\kappa-1)$ then $\mathbb{E}[S]=\pi(h^{2})$ and also $\mathbb{E}[H_{1}H_{2}]=\mathbb{E}[H_{1}]\mathbb{E}[H_{2}]=\pi(h)^{2}$. By Minkowski’s inequality, for $p\geq 1$:
@@
and the terms on the right-hand side are finite by Proposition 43 if $\frac{1}{p}>\frac{2}{m}+\frac{1}{\kappa}$. ∎
Theorem 53.
Under Assumption 2, let $h\in L^{m}(\pi)$ for some $m>2\kappa/(\kappa-2)$. Let $V_{R}$ be as in Definition 51 with $R\geq 0$. For $\pi$-almost all $y$:
@@
By Lemma 52, $\mathbb{E}[S-H_{1}H_{2}]=v(\pi,h)$ if $m>2\kappa/(\kappa-2)$ and $\mathbb{E}\left[\left|S-H_{1}H_{2}\right|^{p}\right]<\infty$ for the range of $p$ given. By Proposition 50, for $\pi$-almost all $y$, $\mathbb{E}[\Phi_{R}]=\pi(h_{0}\cdot g_{y})$ if $m>2\kappa/(\kappa-2)$ and $\mathbb{E}[\left|\Phi_{R}\right|^{p}]$ is finite for the range of $p$ given. Hence, for $m>2\kappa/(\kappa-2)$ we have $\mathbb{E}[V_{R}]=v(P,h)$ and we may conclude the finiteness of $\mathbb{E}\left[\left|V_{R}\right|^{p}\right]$ by Minkowski’s inequality. ∎
Remark 54.
By Minkowski’s inequality we may similarly conclude that any average of estimators of the form given in Definition 51 also has lack-of-bias and moments implied by Theorem 53, and hence that the results apply to (4.4) when $\xi_{k}^{(j)}=1/N^{(j)}$ for $k\in\{1,\ldots,N^{(j)}\}$.
@@ B.7 Ergodic Poisson asymptotic variance estimator
Proposition 55.
Under Assumption 2, let $X$ be a Markov chain with Markov kernel $P$, and $h\in L^{m}(\pi)$ with $m>2\kappa/(\kappa-1)$. For $\pi$-almost all $X_{0}$, the CLT holds for $h$ and for $\pi$-almost all $y$, $v(P,h)=-v(\pi,h)+2\pi(h_{0}\cdot g_{y})$. The estimator (4.2) with $G=G_{y}$, satisfies $\hat{v}(P,h)\to_{{\rm a.s.}}v(P,h)$ as $t\to\infty$.
@@
The assumptions guarantee by Theorem 3 that for $\pi$-almost all $y$, $h\cdot g_{y}\in L^{1}(\pi)$ and since $\mathbb{E}\left[G_{y}(x)\right]=g_{y}(x)$ for $\pi$-almost all $x$ by Theorem 37, the first term on the right-hand side converges almost surely to $\pi(h\cdot g_{y})$ by the Markov chain law of large numbers, while similarly for the second term we have $\pi^{{\rm MC}}(h)\to_{{\rm a.s.}}\pi(h)$ and $\frac{1}{t}\sum_{s=0}^{t-1}G_{y}(X_{s})\to_{{\rm a.s.}}\pi(g_{y})$, so the second term converges almost surely to $\pi(h)\pi(g_{y})$. Hence, the left-hand side converges almost surely to $\pi(h_{0}\cdot g_{y})$, and we conclude. ∎
Theorem 56.
@@
Define $f=(f_{1},\ldots,f_{4})$ with $f_{1}:(x,u)\mapsto h(x)$, $f_{2}:(x,u)\mapsto h(x)^{2}$, $f_{3}=f_{1}\cdot g_{y}$ and $f_{4}=g_{y}$. We observe that
$\hat{v}_{t}(P,h)$ $=\check{\pi}_{t}(f_{1})^{2}-\check{\pi}_{t}(f_{2})+2\check{\pi}_{t}(f_{3})-2\check{\pi}_{t}(f_{1})\check{\pi}_{t}(f_{4}),$
where $\check{\pi}_{t}=\frac{1}{t}\sum_{i=0}^{t-1}\delta_{(X_{i},U_{i})}$ is the empirical measure of the Markov chain introduced in Section B.5. Hence, we define
@@
and $\mu_{Z}=\check{\pi}(f)=(\pi(h),\pi(h^{2}),\pi(h\cdot g_{y}),\pi(g_{y}))$. For each $i\in\{1,\ldots,4\}$, we need to check that $f_{i}\in L^{s}(\check{\pi})$ for appropriately large $s$. We find $f_{1}\in L^{s}(\check{\pi})$ for $s\leq m$, $f_{2}\in L^{s}(\check{\pi})$ for $s\leq m/2$, $f_{3}\in L^{s}(\check{\pi})$ for $s<(\frac{2}{m}+\frac{1}{\kappa})^{-1}$ by Lemma 48 and $f_{4}\in L^{s}(\pi)$ for $s<(\frac{1}{m}+\frac{1}{\kappa})^{-1}$ by Lemma 47. It follows that all of these are in $L^{s}(\check{\pi})$ if $s<(\frac{2}{m}+\frac{1}{\kappa})^{-1}$. The CLT then holds for all $f_{i}$ individually, i.e. there exists $\sigma_{i}^{2}<\infty$ such that
@@
$\sqrt{t}(\hat{v}_{t}(P,h)-v(P,h))=\sqrt{n}(\phi(Z_{n})-\phi(\mu_{Z}))\to_{d}\text{Normal}(0,\sigma^{2}),$
@@
Appendix C Unbiased MCMC estimators
C.1 Unbiased estimators of expectations and signed measures
$H_{k}=h(X_{k})+\sum_{j=1}^{\infty}\{h(X_{k+jL})-h(Y_{k+(j-1)L})\}.$ (C.1)
$H_{k:\ell}$ $=\frac{1}{\ell-k+1}\sum_{t=k}^{\ell}H_{t}$
$v_{t}=|\{s\in\{k,\ldots,\ell\}:\exists j\geq 1:s+jL=t\}|.$
We can re-write this as
$v_{t}=|\{n\in\{t-\ell,\ldots,t-k\}:\frac{n}{L}\in\mathbb{Z}^{+}\}|,$
$v_{t}=\lfloor(t-k)/L\rfloor-\lceil\max(L,t-\ell)/L\rceil+1.$ (C.2)
$H_{k:\ell}=\frac{1}{\ell-k+1}\sum_{t=k}^{\ell}h(X_{t})+\text{BC}_{k:\ell},$ (C.3)
where $\text{BC}_{k:\ell}$ refers to a “bias cancellation” term,
$\text{BC}_{k:\ell}=\sum_{t=k+L}^{\tau-1}\frac{v_{t}}{\ell-k+1}\left\{h(X_{t})-h(Y_{t-L})\right\}.$ (C.4)
$\hat{\pi}({\rm d}x)=\frac{1}{\ell-k+1}\sum_{t=k}^{\ell}\delta_{X_{t}}({\rm d}x)+\sum_{t=k+L}^{\tau-1}\frac{v_{t}}{\ell-k+1}\left\{\delta_{X_{t}}-\delta_{Y_{t-L}}\right\}({\rm d}x),$ (C.5)
as an unbiased approximation of $\pi$. This is of the form
$\hat{\pi}({\rm d}x)=\sum_{n=1}^{N}\omega_{n}\delta_{Z_{n}}({\rm d}x),$ (C.6)
where $N=\max(0,\tau-(k+L))+(\ell-k+1)$, $Z_{n}$ are states from either $(X_{t})$ or $(Y_{t})$ and $\omega_{t}$ are either $(\ell-k+1)^{-1}$, or of the form ±$\pm v_{t}(\ell-k+1)^{-1}$; in particular the weights can be negative.
If we count the cost of sampling from the kernel $P$ as one unit, and the cost of sampling from $\bar{P}$ as one unit if the chains have met already, and two units otherwise, then the random cost of obtaining (C.3), or (C.5), equals $\max(L,\ell+L-\tau)+2(\tau-L)$ units.
C.2 Upper bounds on the distance to stationarity
$|\pi_{t}-\pi|_{\text{TV}}\leq\mathbb{E}\left[\max\left(0,\left\lceil\frac{\tau-L-t}{L}\right\rceil\right)\right].$ (C.7)
Appendix D Comparison with long-run variance estimators
@@
$\hat{\gamma}(u)=\dfrac{1}{m}\sum_{k=1}^{m}\dfrac{1}{t}\sum_{s=0}^{t-u-1}(h(X_{k,s})-\bar{\pi}^{\text{MC}}(h))(h(X_{k,s+u})-\bar{\pi}^{\text{MC}}(h)),$ (D.1)
@@
$v^{\text{sv}}_{b_{t}}(P,h):=\sum_{u=-(t-2)}^{t-2}w\left(\frac{u}{b_{t}}\right)\hat{\gamma}(u)\,.$ (D.2)
@@
$v^{\text{BM}}_{b_{t}}(P,h):=\dfrac{b_{t}}{ma_{t}-1}\sum_{k=1}^{m}\sum_{s=0}^{a_{t}-1}\left(\bar{h}_{k,s}-\bar{\pi}^{\text{MC}}(h)\right)^{2}.$ (D.3)
@@
$v^{\circ}_{r}(P,h):=2v^{\circ}_{b_{t}}(P,h)-v^{\circ}_{\lfloor b_{t}/r\rfloor}(P,h)\,,$ (D.4)
@@
Theorem 57 (Kuelbs & Philipp (1980), Vats et al. (2018)).
Suppose $\mathbb{E}_{\pi}[|h(X)|^{2+\delta}]<\infty$ for some $\delta>0$, and let $P$ be polynomially ergodic of order $\xi>(q+1+\epsilon)/(1+2/\delta)$ for some $q\geq 1$ and $\epsilon>0$. Then, without changing its distribution, process $\{S_{1,t}\}_{t\geq 0}$ can be redefined on a richer probability space together with a standard Brownian motion, $\{B(t)\}_{t\geq 0}$ such that there exists $\sigma>0$, $\lambda>0$, and a finite random variable $D$, such that
@@
Theorem 57 yields a strong invariance principle result under mild conditions on the process, with $\sigma^{2}=v(P,h)$. The existence of a strong invariance principle is one of the sufficient conditions for obtaining the variance of SV and BM estimators. For $q\geq 1$, define
$\Phi^{(q)}=\sum_{u=-\infty}^{\infty}|u|^{q}{\rm cov}_{\pi}\left(h(X_{1,0}),h(X_{1,u})\right)\,.$ (D.5)
@@
Theorem 58 (Bias and variance of SV estimators (Agarwal & Vats 2022, Damerdji 1995, Liu et al. 2022)).
Suppose Theorem 57 holds with $q$ such that
@@
and $b_{t}^{q+1}/t\to 0$ as $t\to\infty$. Then, with $\Phi^{(q)}$ as in (D.5),
$\lim_{t\to\infty}b_{t}^{q}\mathbb{E}\left(v^{\text{sv}}_{r}(P,h)-v(P,h)\right)=-(2-r^{q})k_{q}\Phi^{(q)}\,.$ (D.6)
Additionally, if $\mathbb{E}(D^{4})<\infty$, and $\mathbb{E}_{\pi}(h(X_{1})^{4})<\infty$, then
$\lim_{t\to\infty}\dfrac{t}{b_{t}}{\rm var}\left(v^{\text{sv}}_{r}(P,h)\right)=\dfrac{2}{m}v(P,h)^{2}\int_{-\infty}^{\infty}w_{r}(x)^{2}{\rm d}x\,,$ (D.7)
@@
The bias witnessed in any implementation then depends on the choice of $r$. For $r=1$ (regular SV) the bias is significantly negative, whereas higher choices of $r$ remove the negative bias. Vats & Flegal (2022) recommend going up to $r=3$ to control the effects on the variance of the estimator. Although the rate on the bias depends on the weight function chosen, the rate on the variance does not depend on the weight function. Theorem 58 implies that the variance of the SV estimator vanishes at $t/b_{t}$ rate. In the context of MCMC, the asymptotic distribution of SV estimators have not been studied yet. The bias and variance of BM estimators are similar to SV for $q=1$.
Theorem 59 (Bias and variance of BM estimators (Flegal & Jones 2010, Gupta & Vats 2020, Vats & Flegal 2022)).
Suppose Theorem 57 holds with $q=1$, then
$\lim_{t\to\infty}b_{t}\mathbb{E}\left(v^{\text{BM}}_{r}(P,h)-v(P,h)\right)=-(2-r)\Phi^{(1)}\,.$ (D.8)
Additionally, if $\mathbb{E}(D^{4})<\infty$, and $\mathbb{E}_{\pi}(h(X_{1})^{4})<\infty$, then
$\lim_{t\to\infty}a_{t}{\rm var}\left(v^{\text{BM}}_{r}(P,h)\right)=\dfrac{2}{m}\left[\dfrac{1}{r}+\dfrac{4(r-1)}{r}\right]v(P,h)^{2}\,.$ (D.9)
@@
For the BM estimators as well, the variance has an $a_{t}$ rate of convergence. Given Theorem 59, Damerdji (1995), Flegal & Jones (2010) argue that the mean-square optimal choice of $b_{t}$ is $b_{t}\propto t^{1/3}$, where the proportionality constant may be estimated using parametric techniques discussed in Liu et al. (2022). Recently, Chakraborty et al. (2022) obtained asymptotic normality of single-chain BM estimators $(m=1)$ for geometrically ergodic, reversible chains.
Theorem 60 (Chakraborty et al. (2022)).
Suppose $P$ is geometrically ergodic and $\pi$-reversible, with $\mathbb{E}_{\pi}[h^{8}]<\infty$. If $b_{t}/t^{1/3}\to\infty$, then,
@@
Appendix E Verifying the assumptions in the AR(1) case

### v4

@@ B.6 Unbiased Poisson asymptotic variance estimator
Definition B.28.
Let $H_{1}$ and $H_{2}$ be two independent unbiased estimators of $\pi(h)$ and $S$ be an unbiased estimator of $\pi(h^{2})$, all of the type described in Proposition B.20. Let $R\in\{0,1,\ldots\}$, $\Phi_{R}$ be as in Definition B.26 and define
@@
Lemma B.29.
Under Assumption 2, let $h\in L^{m}(\pi)$ for some $m>2\kappa/(\kappa-1)$, and ${\rm d}\pi_{0}/{\rm d}\pi\leq M$. Let $H_{1}$, $H_{2}$ and $S$ be as in Definition B.28. Then
@@
We note that if $h\in L^{m}(\pi)$ then $h^{2}\in L^{m/2}(\pi)$. Hence, we deduce by Proposition B.20 that if $m/2>\kappa/(\kappa-1)$, i.e. $m>2\kappa/(\kappa-1)$ then $\mathbb{E}[S]=\pi(h^{2})$ and also $\mathbb{E}[H_{1}H_{2}]=\mathbb{E}[H_{1}]\mathbb{E}[H_{2}]=\pi(h)^{2}$. By Minkowski’s inequality, for $p\geq 1$:
@@
and the terms on the right-hand side are finite by Proposition B.20 if $\frac{1}{p}>\frac{2}{m}+\frac{1}{\kappa}$. ∎
Theorem B.30.
Under Assumption 2, let $h\in L^{m}(\pi)$ for some $m>2\kappa/(\kappa-2)$, and ${\rm d}\pi_{0}/{\rm d}\pi\leq M$. Let $V_{R}$ be as in Definition B.28 with $R\geq 0$. For $\pi$-almost all $y$:
@@
By Lemma B.29, $\mathbb{E}[S-H_{1}H_{2}]=v(\pi,h)$ if $m>2\kappa/(\kappa-2)$ and $\mathbb{E}\left[\left|S-H_{1}H_{2}\right|^{p}\right]<\infty$ for the range of $p$ given. By Proposition B.27, for $\pi$-almost all $y$, $\mathbb{E}[\Phi_{R}]=\pi(h_{0}\cdot g_{y})$ if $m>2\kappa/(\kappa-2)$ and $\mathbb{E}[\left|\Phi_{R}\right|^{p}]$ is finite for the range of $p$ given. Hence, for $m>2\kappa/(\kappa-2)$ we have $\mathbb{E}[V_{R}]=v(P,h)$ and we may conclude the finiteness of $\mathbb{E}\left[\left|V_{R}\right|^{p}\right]$ by Minkowski’s inequality. ∎
Remark B.31.
By Minkowski’s inequality we may similarly conclude that any average of estimators of the form given in Definition B.28 also has lack-of-bias and moments implied by Theorem B.30, and hence that the results apply to (4.4) when $\xi_{n}^{(j)}=1/N^{(j)}$ for $n\in\{1,\ldots,N^{(j)}\}$.
@@ B.7 Ergodic Poisson asymptotic variance estimator
Proposition B.32.
Under Assumption 2, let $X$ be a Markov chain with Markov kernel $P$, and $h\in L^{m}(\pi)$ with $m>2\kappa/(\kappa-1)$. For $\pi$-almost all $X_{0}$, the CLT holds for $h$ and for $\pi$-almost all $y$, $v(P,h)=-v(\pi,h)+2\pi(h_{0}\cdot g_{y})$. The estimator (4.2) with $G=G_{y}$, satisfies $\hat{v}_{E}(P,h)\to_{{\rm a.s.}}v(P,h)$ as $t\to\infty$.
@@
The assumptions guarantee by Theorem 3 that for $\pi$-almost all $y$, $h\cdot g_{y}\in L^{1}(\pi)$ and since $\mathbb{E}\left[G_{y}(x)\right]=g_{y}(x)$ for $\pi$-almost all $x$ by Theorem B.14, the first term on the right-hand side converges almost surely to $\pi(h\cdot g_{y})$ by the Markov chain law of large numbers, while similarly for the second term we have $\pi^{{\rm MC}}(h)\to_{{\rm a.s.}}\pi(h)$ and $\frac{1}{t}\sum_{s=0}^{t-1}G_{y}(X_{s})\to_{{\rm a.s.}}\pi(g_{y})$, so the second term converges almost surely to $\pi(h)\pi(g_{y})$. Hence, the left-hand side converges almost surely to $\pi(h_{0}\cdot g_{y})$, and we conclude. ∎
Theorem B.33.
@@
Recall the notation of Appendix B.5 with $\bar{g}_{y}$ and $\check{\pi}$. Define $f=(f_{1},\ldots,f_{4})$ with $f_{1}:(x,u)\mapsto h(x)$, $f_{2}:(x,u)\mapsto h(x)^{2}$, $f_{3}=f_{1}\cdot\bar{g}_{y}$ and $f_{4}=\bar{g}_{y}$. We observe that (4.2) can be rewritten as
$\hat{v}_{E}(P,h)$ $=\check{\pi}_{t}(f_{1})^{2}-\check{\pi}_{t}(f_{2})+2\check{\pi}_{t}(f_{3})-2\check{\pi}_{t}(f_{1})\check{\pi}_{t}(f_{4}),$
where $\check{\pi}_{t}=\frac{1}{t}\sum_{i=0}^{t-1}\delta_{(X_{i},U_{i})}$ is the empirical measure of the Markov chain with transition $T$ introduced in Section B.5. Hence, we define
@@
and $\mu_{Z}=\check{\pi}(f)=(\pi(h),\pi(h^{2}),\pi(h\cdot g_{y}),\pi(g_{y}))$. For each $i\in\{1,\ldots,4\}$, we need to check that $f_{i}\in L^{s}(\check{\pi})$ for appropriately large $s$. We find $f_{1}\in L^{s}(\check{\pi})$ for $s\leq m$, $f_{2}\in L^{s}(\check{\pi})$ for $s\leq m/2$, $f_{3}\in L^{s}(\check{\pi})$ for $s<(\frac{2}{m}+\frac{1}{\kappa})^{-1}$ by Lemma B.25 and $f_{4}\in L^{s}(\pi)$ for $s<(\frac{1}{m}+\frac{1}{\kappa})^{-1}$ by Lemma B.24. It follows that all of these are in $L^{s}(\check{\pi})$ if $s<(\frac{2}{m}+\frac{1}{\kappa})^{-1}$. The CLT then holds for all $f_{i}$ individually, i.e. there exists $\sigma_{i}^{2}<\infty$ such that
@@
$\sqrt{t}(\hat{v}_{t}(P,h)-v(P,h))=\sqrt{t}(\phi(Z_{t})-\phi(\mu_{Z}))\to_{d}\text{Normal}(0,\sigma^{2}),$
@@
Appendix C Comparison with long-run variance estimators
@@
$\hat{\gamma}(u)=\dfrac{1}{m}\sum_{k=1}^{m}\dfrac{1}{t}\sum_{s=0}^{t-u-1}(h(X_{k,s})-\bar{\pi}^{\text{MC}}(h))(h(X_{k,s+u})-\bar{\pi}^{\text{MC}}(h)),$ (C.1)
@@
$v^{\text{sv}}_{b_{t}}(P,h):=\sum_{u=-(t-2)}^{t-2}w\left(\frac{u}{b_{t}}\right)\hat{\gamma}(u)\,.$ (C.2)
@@
$v^{\text{BM}}_{b_{t}}(P,h):=\dfrac{b_{t}}{ma_{t}-1}\sum_{k=1}^{m}\sum_{s=0}^{a_{t}-1}\left(\bar{h}_{k,s}-\bar{\pi}^{\text{MC}}(h)\right)^{2}.$ (C.3)
@@
$v^{\circ}_{r}(P,h):=2v^{\circ}_{b_{t}}(P,h)-v^{\circ}_{\lfloor b_{t}/r\rfloor}(P,h)\,,$ (C.4)
@@
Theorem C.34 (Kuelbs & Philipp (1980), Vats et al. (2018)).
Suppose $\mathbb{E}_{\pi}[|h(X)|^{2+\delta}]<\infty$ for some $\delta>0$, and let $P$ be polynomially ergodic of order $\xi>(q+1+\epsilon)/(1+2/\delta)$ for some $q\geq 1$ and $\epsilon>0$. Then, without changing its distribution, process $\{S_{1,t}\}_{t\geq 0}$ can be redefined on a richer probability space together with a standard Brownian motion, $\{B(t)\}_{t\geq 0}$ such that there exist $\sigma>0$, $\lambda>0$, and a finite random variable $D$, such that
@@
Theorem C.34 yields a strong invariance principle result under mild conditions on the process, with $\sigma^{2}=v(P,h)$. The existence of a strong invariance principle is one of the sufficient conditions for obtaining the variance of SV and BM estimators. For $q\geq 1$, define
$\Phi^{(q)}=\sum_{u=-\infty}^{\infty}|u|^{q}{\rm cov}_{\pi}\left(h(X_{1,0}),h(X_{1,u})\right)\,.$ (C.5)
@@
Theorem C.35 (Bias and variance of SV estimators (Agarwal & Vats 2022, Damerdji 1995, Liu et al. 2022)).
Suppose Theorem C.34 holds with $q$ such that
@@
and $b_{t}^{q+1}/t\to 0$ as $t\to\infty$. Then, with $\Phi^{(q)}$ as in (C.5),
$\lim_{t\to\infty}b_{t}^{q}\mathbb{E}\left(v^{\text{sv}}_{r}(P,h)-v(P,h)\right)=-(2-r^{q})k_{q}\Phi^{(q)}\,.$ (C.6)
Additionally, if $\mathbb{E}(D^{4})<\infty$, and $\mathbb{E}_{\pi}[h(X)^{4}]<\infty$, then
$\lim_{t\to\infty}\dfrac{t}{b_{t}}{\rm var}\left(v^{\text{sv}}_{r}(P,h)\right)=\dfrac{2}{m}v(P,h)^{2}\int_{-\infty}^{\infty}w_{r}(x)^{2}{\rm d}x\,,$ (C.7)
@@
The bias witnessed in any implementation then depends on the choice of $r$. For $r=1$ (regular SV) the bias is significantly negative, whereas higher choices of $r$ remove the negative bias. Vats & Flegal (2022) recommend going up to $r=3$ to control the effects on the variance of the estimator. Although the rate on the bias depends on the weight function chosen, the rate on the variance does not depend on the weight function. Theorem C.35 implies that the variance of the SV estimator vanishes at $t/b_{t}$ rate. In the context of MCMC, the asymptotic distribution of SV estimators have not been studied yet. The bias and variance of BM estimators are similar to SV for $q=1$.
Theorem C.36 (Bias and variance of BM estimators (Flegal & Jones 2010, Gupta & Vats 2020, Vats & Flegal 2022)).
Suppose Theorem C.34 holds with $q=1$, then
$\lim_{t\to\infty}b_{t}\mathbb{E}\left(v^{\text{BM}}_{r}(P,h)-v(P,h)\right)=-(2-r)\Phi^{(1)}\,.$ (C.8)
Additionally, if $\mathbb{E}(D^{4})<\infty$, and $\mathbb{E}_{\pi}[h(X)^{4}]<\infty$, then
$\lim_{t\to\infty}a_{t}{\rm var}\left(v^{\text{BM}}_{r}(P,h)\right)=\dfrac{2}{m}\left[\dfrac{1}{r}+\dfrac{4(r-1)}{r}\right]v(P,h)^{2}\,.$ (C.9)
@@
For the BM estimators as well, the variance has an $a_{t}$ rate of convergence. Given Theorem C.36, Damerdji (1995), Flegal & Jones (2010) argue that the mean-square optimal choice of $b_{t}$ is $b_{t}\propto t^{1/3}$, where the proportionality constant may be estimated using parametric techniques discussed in Liu et al. (2022). Recently, Chakraborty et al. (2022) obtained asymptotic normality of single-chain BM estimators $(m=1)$ for geometrically ergodic, reversible chains.
Theorem C.37 (Chakraborty et al. (2022)).
Suppose $P$ is geometrically ergodic and $\pi$-reversible, with $\mathbb{E}_{\pi}[h(X)^{8}]<\infty$. If $b_{t}/t^{1/3}\to\infty$, then,
@@
Appendix D Verifying the assumptions in the AR(1) case